how do i do 40% of x is 28 what is x. i dont get it
Answer by Alan3354(69443)

how do i do 40% of x is 28 what is x. i dont get it
40% is 40/100 = 0.4
---------------------
40% of x = 0.4x = 28
x = 28/0.4 = 280/4
x = 70